OpenStudy (anonymous):

let v=(2,8) and w=(-1,-4). which of the following is true A. v=-2w B. The x-component of v is 2 C.V*W=-32 D. V and W are perpendicular

OpenStudy (anonymous):

let's check one by one: A. v = -2w -2w = -2(-1,-4) = (-2*(-1),-2*(-4)) = (2, 8)

OpenStudy (anonymous):

so, v = -2w is True

OpenStudy (anonymous):

and b is? how would you do that?

OpenStudy (anonymous):

V = (2, 8) is a vector with x-componet 2 and y-component 8

OpenStudy (anonymous):

v*w = (2,8)*(-1,-4) = 2*(-1) + 8*(-4) = -2 - 32 = -34

OpenStudy (anonymous):

so B is false?

OpenStudy (anonymous):

B is true, since v's x-componet is 2

OpenStudy (anonymous):

oh okay i see

OpenStudy (anonymous):

c is false, v*w = -34 and not -32

OpenStudy (anonymous):

as v*w is not equal to zero, v and w are not perpendicular

OpenStudy (anonymous):

So only A and B are true, thanks

OpenStudy (anonymous):

that's it

OpenStudy (anonymous):

Np ;)

OpenStudy (anonymous):

I have one more question?:)

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

evaluate the following expression: -1(5,-1)+2(1,1)

OpenStudy (anonymous):

what you have done so far?

OpenStudy (anonymous):

multiply the two out side numbers?

OpenStudy (anonymous):

show your work, then I see where you are getting wrong (if so)

OpenStudy (anonymous):

(-5,1)+(2,2)

OpenStudy (anonymous):

that's correct, now just sum component by component

OpenStudy (anonymous):

(-3,3)?

OpenStudy (anonymous):

You got it!

OpenStudy (anonymous):

Thanks:D

OpenStudy (anonymous):

No problem ^^